OpenStudy (anonymous):

one number is 5 more than twice the other. their sum is 26. let x represent the smaller number someone please help me

OpenStudy (anonymous):

I got x=7.\[7\times2=14\]\[14+5=19\]\[19+7=26\] Therefore, my answer is 7

OpenStudy (anonymous):

okay this a multiple choice question so which answer would ne correct a. x+2+(x+5)=26 b. x+(x+5)=26 c. x+2(x) d. x+(2x+5)=26 e. not given

OpenStudy (anonymous):

Just give me a second. :)

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

In the equation 'D', the parenthesis show the number being multiplied by 2 like I did, and then adding 5. Then, it just adds the original value of x. It's basically exactly what I did. :)

OpenStudy (anonymous):

And, welcome to OpenStudy.

OpenStudy (anonymous):

okay so I have another problem similar to this one and thank you so much

OpenStudy (anonymous):

Ok, you can post it in another question or post it here. I'll help you either way.

OpenStudy (anonymous):

You can tag me to the question.

OpenStudy (anonymous):

which of the following expressions represents the product of 3 less than twice x and 2 more than the quantity 3 times x ? a. -6x^2+25x+6 b. 6x^2+5x+6 c. 6x^2-5x+6 d. 6x^2-5x-6 e. 6x^2-13x-6

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

I don't know, this one is a little more difficult. Sorry I can't help. At least I helped with the other one.